OpenStudy (anonymous):

Simplify the given expression √-64/(7-6i) - (2-2i) a) 64+40i/-3 b) -32+40i/41 c) -64+40i/-3 d) 32+40i/41

OpenStudy (lalaly):

\[\frac{\sqrt{-64}}{7-6i}-(2-2i)\] is this it?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

wait... no the -(2-2i) goes next to 7-6i

OpenStudy (lalaly):

so theyr both in the denominator?

OpenStudy (anonymous):

yes

OpenStudy (lalaly):

\[=\frac{8i}{5-4i}\]\[=\frac{8i}{5-4i} \times \frac{5+4i}{5+4i} = \frac{40i-32}{25-16}\]

OpenStudy (lalaly):

\[=\frac{40i-32}{9}\]